KAPALOTJAY:

can yall help me solve this 6x(9+8-9x)=?

addisonwk:

Use algebra calculator lol that's what i do

MiraAngel:

You need to distribute 6x to all of those different parts in the parentheses. Take 6x times 9. Then take 6x times 8. Then take 6x times -9x. 6x*9= 54x 6x*8= 48x 6x*-9x= -54x^2 So... 54x+48x+ (-54x^2) 102x-54x^2

KAPALOTJAY:

thx

MiraAngel:

I don't know how far you want it reduced. The problem may just want you to reduce it to the point where the problem is solved within the parentheses like this.... 6x(9+8-9x)=6x(-9x+17)

KAPALOTJAY:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MiraAngel I don't know how far you want it reduced. The problem may just want you to reduce it to the point where the problem is solved within the parentheses like this.... 6x(9+8-9x)=6x(-9x+17) \(\color{#0cbb34}{\text{End of Quote}}\) thx